# Strategic Learning and Trading in Broker-Mediated Markets

### v1

@@ Appendix A Proof of Proposition 1
$\mathrm{d}\mathbb{V}^{\text{alt},B}_{t}=\Big{\{}(\sigma^{\alpha})^{2}-2\,\kappa^{\alpha}\,\mathbb{V}^{\text{alt},B}_{t}-(r_{t}^{M})^{2}\Big{\}}\,\mathrm{d}t\,.$ (A.12)
[…]
$\mathrm{d}\left(\epsilon\,\alpha\right)_{t}$ $=\epsilon_{t}\,\mathrm{d}\alpha_{t}+\alpha_{t}\,\mathrm{d}\epsilon_{t}+\mathrm{d}\langle\epsilon,\alpha\rangle_{t}$ (A.13)
$=-\kappa^{\alpha}\,\epsilon_{t}\,\alpha_{t}\,\mathrm{d}t+\sigma^{\alpha}\,\epsilon_{t}\,\mathrm{d}W_{t}^{\alpha}+i\,\epsilon_{t}\,\alpha_{t}\,r(t)\,\mathrm{d}Z_{t}+i\,\sigma^{\alpha}\,\epsilon_{t}\,r(t)\,\mathfrak{K}(t)\,\mathrm{d}t\,,$ (A.14)
$\mathrm{d}\left(\epsilon\,\hat{\alpha}^{\text{alt}}\right)_{t}$ $=\epsilon_{t}\,\mathrm{d}\hat{\alpha}^{\text{alt}}_{t}+\hat{\alpha_{t}}^{\text{alt}}\,\mathrm{d}\epsilon_{t}+\mathrm{d}\langle\epsilon,\hat{\alpha}^{\text{alt}}\rangle_{t}$ (A.15)
$=-\kappa^{\alpha}\,\epsilon_{t}\,\hat{\alpha}^{\text{alt}}_{t}\,\mathrm{d}t+\epsilon_{t}\,r^{M}_{t}\,\mathrm{d}I_{t}+i\,\epsilon_{t}\,r(t)\,\hat{\alpha}^{\text{alt}}_{t}\,\mathrm{d}Z_{t}+i\,\epsilon_{t}\,r(t)\,r_{t}^{M}\,\mathrm{d}t\,.$ (A.16)
Note that by applying the tower property, $\mathbb{E}\left[\epsilon_{t}(\hat{\alpha}^{\text{alt}}_{t}-\alpha_{t})\right]=0$. By subtracting (A.14) from (A.16) and taking the total expectation, we have
[…]
We know that $\mathbb{V}^{\text{alt},B}_{t}\leq\text{Var}(\alpha_{t})=\frac{(\sigma^{\alpha})^{2}}{2\,\kappa^{\alpha}}\,\exp\left(-2\,\kappa^{\alpha}\,t\right)$, so that $\mathbb{E}\left[\int_{0}^{T}\left(\hat{\alpha}^{\text{alt}}_{u}-\alpha_{u}\right)^{2}\,\mathrm{d}u\right]<\infty$. Thus, the rightmost term in (A.17) equals zero. We obtain

### v2

@@ B.2  Proof of Proposition 1
$\mathrm{d}\mathbb{V}^{\text{alt},B}_{t}=\Big\{(\sigma^{\alpha})^{2}-2\,\kappa^{\alpha}\,\mathbb{V}^{\text{alt},B}_{t}-(r_{t}^{M})^{2}\Big\}\,\mathrm{d}t\,.$ (B.30)
Now, we wish to determine the process $r^{M}$. Define $\epsilon_{t}:=\exp\left(i\,\int_{0}^{t}r(u)\,\mathrm{d}Z_{u}+\frac{1}{2}\int_{0}^{t}|r(u)|^{2}\,du\right)$ where $r:[0,T]\to\mathbb{R}$ is measurable and bounded. By Ito’s formula, $\mathrm{d}\epsilon_{t}=i\,\epsilon_{t}\,r(t)\,\mathrm{d}Z_{t}$. We calculate
$\mathrm{d}\left(\epsilon\,\alpha\right)_{t}$ $=\epsilon_{t}\,\mathrm{d}\alpha_{t}+\alpha_{t}\,\mathrm{d}\epsilon_{t}+\mathrm{d}\langle\epsilon,\alpha\rangle_{t}$ (B.31)
$=-\kappa^{\alpha}\,\epsilon_{t}\,\alpha_{t}\,\mathrm{d}t+\sigma^{\alpha}\,\epsilon_{t}\,\mathrm{d}W_{t}^{\alpha}+i\,\epsilon_{t}\,\alpha_{t}\,r(t)\,\mathrm{d}Z_{t}+i\,\sigma^{\alpha}\,\epsilon_{t}\,r(t)\,\mathfrak{K}(t)\,\mathrm{d}t\,$, (B.32)
$\mathrm{d}\left(\epsilon\,\hat{\alpha}^{\text{alt}}\right)_{t}$ $=\epsilon_{t}\,\mathrm{d}\hat{\alpha}^{\text{alt}}_{t}+\hat{\alpha_{t}}^{\text{alt}}\,\mathrm{d}\epsilon_{t}+\mathrm{d}\langle\epsilon,\hat{\alpha}^{\text{alt}}\rangle_{t}$ (B.33)
$=-\kappa^{\alpha}\,\epsilon_{t}\,\hat{\alpha}^{\text{alt}}_{t}\,\mathrm{d}t+\epsilon_{t}\,r^{M}_{t}\,\mathrm{d}I_{t}+i\,\epsilon_{t}\,r(t)\,\hat{\alpha}^{\text{alt}}_{t}\,\mathrm{d}Z_{t}+i\,\epsilon_{t}\,r(t)\,r_{t}^{M}\,\mathrm{d}t\,$, (B.34)
where
Note that by applying the tower property, $\mathbb{E}\left[\epsilon_{t}(\hat{\alpha}^{\text{alt}}_{t}-\alpha_{t})\right]=0$. By subtracting (B.32) from (B.34) and taking the total expectation, we have
$\begin{split}0&=\mathbb{E}\left[\int_{0}^{t}i\,\epsilon_{u}\,r(u)\,\left(\hat{\alpha}^{\text{alt}}_{u}-\alpha_{u}\right)\,\mathrm{d}Z_{u}\right]+\int_{0}^{t}i\,r(u)\,\mathbb{E}\left[\epsilon_{u}\,(r_{u}^{M}-\sigma^{\alpha}\,\mathfrak{K}(u))\right]\,\mathrm{d}u\\ &=\int_{0}^{t}i\,r(u)\,\mathbb{E}\left[\epsilon_{u}\,\left\{r_{u}^{M}-\sigma^{\alpha}\,\mathfrak{K}(u)+(\hat{\alpha}^{\text{alt}}_{u}-\alpha_{u})\,\mathfrak{G}_{7}(u)\,\alpha_{u}\right\}\right]\,\mathrm{d}u+\mathbb{E}\left[\int_{0}^{t}i\,\epsilon_{u}\,r(u)\,\left(\hat{\alpha}^{\text{alt}}_{u}-\alpha_{u}\right)\,\mathrm{d}\tilde{W}_{u}\right]\,.\end{split}$ (B.35)
We know that $\mathbb{V}^{\text{alt},B}_{t}\leq\text{Var}(\alpha_{t})=\frac{(\sigma^{\alpha})^{2}}{2\,\kappa^{\alpha}}\,\exp\left(-2\,\kappa^{\alpha}\,t\right)$, so that $\mathbb{E}\left[\int_{0}^{T}\left(\hat{\alpha}^{\text{alt}}_{u}-\alpha_{u}\right)^{2}\,\mathrm{d}u\right]<\infty$. Thus, the rightmost term in (B.35) equals zero. We obtain